The rplain package

Axel Sommerfeldt axel@ang-physik.uni-kiel.de

1994/10/21

Abstract The rplain package redefines the ‘plain’ pagestyle. The page numbers

are now in the lower right corner instead of being centered, like in this onesided document. In twosided documents, the page numbers are in the lower left or right corner.

1 The user interface

To use this package just type

   \usepackage{rplain}

in the preamble of your document.

2 The code

2.1 Identification

\NeedsTeXFormat{LaTeX2e}[1994/06/01]
\ProvidesPackage{rplain}[1994/10/10 v1.0 rplain package (AS)]
\typeout{Package: rplain v1.0 <1994/10/10> (Axel Sommerfeldt)}

2.2 Main code

\ps@plain This macro was taken from report.cls (LaTeX2e 94/06/01 patch level 3) and

modified.

\renewcommand{\ps@plain}{%
   \renewcommand{\@mkboth}{\@gobbletwo}%
   \renewcommand{\@oddhead}{}%
   \renewcommand{\@evenhead}{}%
   \renewcommand{\@evenfoot}{\reset@font\rmfamily\thepage\hfil}%
   \renewcommand{\@oddfoot}{\reset@font\rmfamily\hfil\thepage}}

Now we have to call \pagestyle{plain} (again), so that the above macro will be used.

\pagestyle{plain}